A speed boat can travel at 16km/h in still water. A 60km course takes 2 hours longer against the current than with it. Find the speed of the current.
Answer by mananth(16946)

let x be the rate of current
downstream speed = (16+x)
upstream speed = (16-x)
t=d/r
time upstream - time downstream = 2 hours
distance = 60

60/(16-x) -60/(16+x) =2
60(16+x) - 60(16-x)=2(16+x)(16-x)

960+60x -960+60x=2(256-x^2)
120x+2x^2-512=0
2x^2+120x-512=0
/2
x^2+60x-256=0
x^2+64x-4x-256=0
x(x+64)-4(x+64)=0
(x-4)(x+64)=0
x=4 the positive value